Hi, Can you please help me
2/y + 1 = 17
Thanks heaps

Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

2/y=17-1
2/y=16
2=16y
2/16=y
y=1/8

Answer by stanbon(75887)

2/y + 1 = 17
----
Subtract 1 from each side:
2/y = 16
----
Divide each side by 2
1/y = 8
---
Cross-multiply
8y = 1
---
Divide both sides by 8
y = 1/8
-----------------
cheers,
Stan H.